OpenStudy (saylilbaby):

let f(x)=x^2+7 andg(x)=x+3/x. Find (g*f)(-5)

OpenStudy (saylilbaby):

@aaronq @confluxepic @ganeshie8 @Luigi0210 @hartnn @Luigi0210

OpenStudy (anonymous):

Plug in -5 for the x in f(x). Do the same for g(x). Multiply the two results

OpenStudy (serenity74):

(-5)^2 + 7 = (answer) g(x) (answer)+ 3 / (answer) basically what @peachpi said

OpenStudy (saylilbaby):

im stuck @Serenity74 did you get 32/35 or 35/12

OpenStudy (serenity74):

Didn't calculate it wait a sec.

OpenStudy (saylilbaby):

k @Serenity74

OpenStudy (serenity74):

I didnt get 32/35

OpenStudy (saylilbaby):

what you got @Serenity74

OpenStudy (serenity74):

25 + 7 = 32 (x = 32) 32 + 3 = 35 (x +3) Then it would be 35/x So it would be 35/32

OpenStudy (michele_laino):

hint: by definition of function of function, we can write this: \[g\left( {f\left( x \right)} \right) = f\left( x \right) + \frac{3}{{f\left( x \right)}} = \left( {{x^2} + 7} \right) + \frac{3}{{{x^2} + 7}}\] now, please replace \(x=-5\) at the right side

OpenStudy (serenity74):

25 + 7 = 32 (x = 32) 32 + 3 = 35 (x +3) Then it would be 35/x So it would be 35/32

OpenStudy (serenity74):

Do you get or no???

OpenStudy (serenity74):

Do you get or no???

OpenStudy (serenity74):

Do you get or no???

OpenStudy (serenity74):

Do you get or no???

OpenStudy (serenity74):

Do you get or no???

OpenStudy (serenity74):

Do you get or no???

OpenStudy (serenity74):

Do you get or no???